Let 𝑆 be any non-empty set and let 𝑉 (𝑆) be the set of all real valued functions on ℝ. Define addition on 𝑉 (𝑠) by (𝑓 + 𝑔)(π‘₯) = 𝑓 (π‘₯) + 𝑔(π‘₯) and scalar multiplication by (𝛼 β‹… 𝑓 )(π‘₯) = 𝛼𝑓 (π‘₯). Check that (𝑉 (𝑆), +, β‹…) is a vector space.


1
Expert's answer
2022-05-10T23:13:02-0400

Recall the definition of the vector space:

A vector space over "\mathbb{C}" is a set "V" with operations of addition:"V\times V\rightarrow V" and scalar multiplication: "{\mathbb{C}}\times V\rightarrow V" satisfying the following properties:

  1. Commutativity: "f+g=g+f" for all "f,g\in V".
  2. Associativity: "(f+g)+h=f+(g+h)" for all "f,g,h\in V".
  3. Additive identity: there exists "0\in V": "0+f=f" for all "f\in V".
  4. Additive inverse: For every "f\in V" there exists "g\in V" satisfying: "f+g=0".
  5. Multiplicative identity: there is "1f=f" for all "f\in V".
  6. Distributivity: "a(f+g)=af+ag" and "(a+b)f=af+bf" for all "a,b\in{\mathbb{C}}" and all "f\in V".

We will check the properties for "V(S)". Properties "1" and "2" are satisfied due to the properties of standard addition and definition: "(f+g)(x)=f(x)+g(x)". Consider zero function . It satisfies property "3". For any function "f(x)" consider the function "g(x)=-f(x)". It satisfies property "4" . Properties "5" and "6" are satisfied due to the definition "(\alpha\cdot f)(x)=\alpha f(x)" and properties of standard multiplication.